[Brahms-l] Shift report 20050219 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 19 2005 - 00:05:50 EST
Shifter leader was HI
Shifters were HHD


1. Beam 1700-1800 and 1930-2330. Due to a large initial loss of ions in the blue ring the first beam was dumped out of schedule. Apart from this no problems.

2.Runs: 13246,13255,13256,13257 
  FS@4B3450: 114k,266k,235k,115k
  MRS@40A1050: 89k,174k,178k,144k

3. Liquid started leaking out of the magnet power supply cooling system. This should not yield any measurement problems according to the magnet crew.
   


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 19 00:06:21 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 19 2005 - 00:06:34 EST